Are points A (2, 5), B (4,9) and C (6, 15) collinear? show your calculations.
thanks for your help!

Answer by mananth(16946)

Are points A (2, 5), B (4,9) and C (6, 15) collinear? show your calculations.
if the slope of AB = slope of AC then the points are co-linear
..
slope AB = 9-5/4-2
slope Ab = 2
...
Slope AC = 15-5/6-2
slope AC = 5/2
...
Hence they are not co-linear